If a random variable X follows Normal distribution such that P (9,6<X<13.8)=0.07008 and P(X>9.6)= 0.8159. Find the mean & variance of X and find P(X> 13.8/X>9.6).


1
Expert's answer
2020-12-29T14:54:24-0500

We remind that the probability density function for the normal distribution has the form: "p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}" . We have: "\int_{9.6}^{13.8}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}dx=0.07008" and "\int_{9.6}^{+\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}dx=0.8159" . We make the change: "z=\frac{x}{\sigma}-\frac{\mu}{\sigma}=\frac{x}{\sigma}-\alpha" and receive: "\int_{\frac{9.6}{\sigma}-\alpha}^{\frac{13.8}{\sigma}-\alpha}\frac{1}{\sqrt{2\pi}}e^{-\frac12z^2}dz=0.07008" and "\int_{\frac{9.6}{\sigma}-\alpha}^{+\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12z^2}dz=0.8159" with "\alpha=\frac{\sigma}{\mu}" .

From the last integral we receive that: "\frac{9.6}{\sigma}-\alpha=-0.9". We substitute that in the first one and get: "-0.663=\frac{13.8}{\sigma}-\alpha" . We receive: "\frac{4,2}{\sigma}=0,237". Thus, we get: "\sigma=17,722" and "\mu=1.442." "P(X>13.8)=\int_{13.8}^{+\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}dx\approx0.2428" and

"P(X>9.6)=\int_{9.6}^{+\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}dx\approx0.3226".

"P(X>13.8/X>9.6)=\frac{P((X>13.8)\, \cap\, (X>9.6))}{P(X>9.6)}=\frac{P(X>13.8)}{P(X>9.6)}=\frac{0.2428}{0.3226}\approx0.7526"

We used the following code in Anaconda for computations:


from scipy import integrate

import numpy as np

import math


func = lambda x:(1/(17.722*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-1.442)/17.722)*((x-1.442)/17.722))


e = integrate.quad(func, 9.6,+np.inf)

print(e)

We subsituted different values of "\mu" and "\sigma" in it.